# The effect of friction on the dynamics of targeted energy transfer by symmetric vibro-impact dampers

### v1

@@ 6 Conclusion
This study provides a comprehensive analysis of the nonlinear dynamics of a horizontal symmetric vibro-impact nonlinear energy sink with dry friction, extending previous frictionless models. The system with friction has been analyzed using two complementary methods, offering a deeper understanding of the system’s behavior.
A key contribution is the extension of the multiple scales method for vibro-impact NES to explicitly incorporate friction, enabling the analytical derivation of the slow invariant manifold for two-impacts-per-period solutions. This inclusion of frictional terms broadens the applicability of MSM, overcoming previous limitations of frictionless models. To complement the MSM, an impact map approach is introduced to address its limitations in capturing fast dynamics. While the MSM provides an explicit analytical expression of the SIM and a first-order approximation of the system’s slow dynamics on it, the impact map represents the fast evolution of trajectories through discrete impacts. The strong agreement between these methods reinforces the validity of multiple scales analysis and highlights the robustness of the impact map approach in predicting periodic solutions, their stability and bifurcations. As the mass ratio approaches zero, both methods converge, further validating their accuracy at different approximation levels. Another key contribution is the extension of MSM to approximate the frequency response curves for various response regimes while incorporating frictional effects. This reveals the coexistence of two distinct response regimes at the same amplitude level —one dominated by impacts and the other by continuous sliding. The analysis further demonstrates how friction alters the activation threshold of the VI-NES, influences the stability of periodic solutions, and suppresses strongly modulated responses near resonance. Additionally, discontinuities in the FRC due to chaotic windows and fold bifurcations are identified using the impact map technique. A brute-force bifurcation diagram confirmed the analytical findings and the computed branches of periodic solutions. It is shown that higher amplitude levels correspond to periodic windows involving more impacts per cycle.
This work establishes a rigorous framework for understanding the dynamics of VI-NES subjected to friction, and offers analytical and numerical tools for system design and optimization. The findings offer valuable insights for improving the performance of nonlinear energy sinks in practical applications.

### v2

@@ 6 Conclusion
This study presents a comprehensive analytical and numerical investigation of the nonlinear dynamics of a symmetric vibro-impact nonlinear energy sink (VI-NES) with dry friction, extending previously idealized frictionless models toward more realistic configurations by accounting for dry friction. The analysis combines the Multiple Scales Method (MSM) and the impact-map approach within a unified framework that captures both the slow and fast time-scale dynamics of frictional vibro-impact systems.
The first key contribution is the extension of the MSM to include friction, enabling the derivation of closed-form expressions for the slow invariant manifold of the two-impacts-per-period ($2$-IPP) regime and a first-order approximation of the system’s slow dynamics on it. This inclusion broadens the applicability of the MSM and provides deeper insight into the nonlinear near-resonant behavior of the system. It also yields closed-form expressions for design-relevant performance metrics, such as the VI-NES activation threshold, effective dissipation characteristics, and backbone curve. These metrics define the engagement conditions and energy-transfer efficiency of the absorber during targeted energy transfer.
The complementary impact-map approach captures the fast evolution of trajectories through discrete impact events and provides closed-form expressions for the amplitude bounds and stability limits of the $2$-IPP regime. Together, these analytical results define the optimal operating range of the VI-NES and establish a consistent foundation for parameter tuning and design optimization. The strong agreement between both methods reinforces the validity of the multiple scales analysis and demonstrates the robustness of the impact-map approach in predicting periodic responses, their stability, and bifurcations.
Another key contribution is the use of the extended MSM formulation to investigate the coexistence of response regimes near the optimal operating range of the VI-NES. The analysis reveals two distinct responses at the same amplitude levels: one dominated by impacts and the other governed by continuous sliding. Numerical results validate these analytical predictions and the computed branches of periodic solutions, confirming the transitions between regimes as excitation conditions vary. The study also demonstrates that friction introduces competing effects that depend strongly on vibration amplitude. At low amplitudes, friction enhances dissipation and raises the activation threshold, improving energy-transfer efficiency but reducing robustness. At higher amplitudes, however, its influence diminishes, and the system behavior converges toward the idealized frictionless case. This dual role emphasizes the need to balance robustness and efficiency when tuning VI-NES parameters for optimal performance.
This work establishes a rigorous framework for understanding the dynamics of VI-NES subjected to friction and provides analytical and numerical tools for system design and optimization. It also forms the theoretical foundation for a forthcoming experimental study that will validate the analytical results and assess the VI-NES performance under real excitation conditions.